Air at a temperature of 300 K flows over one side of a flat plate of width 1 m at a velocity of 20 m/s. The plate has a constant surface temperature of 350 K. Assume Re(x,c)=5x10^5.

a) What is the velocity boundary layer thickness at the end of the plate if L=0.25 m? What if L=1 m?

b) Calculate the drag on the plate if L=0.25 m. What is the drag if L=1 m?

c) Find the heat transfer rate from the plate to the air if L=0.25 m. What is the heat transfer rate if L=1 m.

d) A small diameter wire is placed at the leading edge of the flat plate. This wire trips the flow, causing it to be turbulent over the entire length of the plate. Repeat the calculations performed in parts a), b), and c) for this condition.
Answer #1

Given: Temperature of flowing air To = 300 K Temperature of plate Tp = 350 K Velocity of air flow of flat plate Vair = 20 m WRez-035 = Vair L - 20x0.25 <= 2.76x10 - = Rea VaivL20x1 V 1.807x10-5 -5 = 11.06x103 At L=0.25 m, Reynolds number is less that0.38L 0.38x1 7 = 0.023 m | = = 0.382 Ref (11.06x10°)* Oza1 = 23 mm (b) Calculate drag on the plate if L=0.25 m and L=1 m. At20x. 5x10ʻ =- 1.807x10-5 x = 0.451 m 7. – 9.451 – 0,451 5, mixed boundary layer condition exists. Re So, for mixed boundary lEp =(Cs)-ox *324,73 Fp = x= 0.0025x2x1.086x(0.25x1)x 20° F) = 0.135 N Drag force for L=1 m length of plate. Fr= C, ).mx5p4,Vä(c) Calculate heat transfer from the plate at L=0.25 m and L=1 m. For plate length L=0.25 m, flow is laminar. Average ConvectFor plate length L=1 m, mixed boundary layer condition exists. Average convective heat transfer coefficient for mixed boundarHeat transfer for plate length L=0.25 m. 22-0.25 = H20.25A,(T. -T.) Ql+0.25 = 34.71(0.25x1)(350 – 300) Ql=0.25 = 433.875 W He0.38L 0.38x0.25 F = 0.0068 m turb ) L=0.25 11-025 Res (5x109) (Sterb )2-0.25 = 6.8 mm Velocity boundary layer at L=1m, Reyno(C:)., = 0.074 Re * = 0.074x(11.06x109)*** = 0.0045 Drag force at x=0.25 m length of plate. (Fo],028 =(,)x xz p4,Vidal (Fo )1r=1 (F) = (Cs)->p4,V (FD),sa = 0.0045*5*1.086x(1x1)x202 (FD)== 0.977 N = Convective heat transfer coefficient for turbulent fha=0.25 +0.25 =1191.11 0.028 hx=0.25 = 133.404 W/ /m²K At x=1 m, Reynolds number will be same as before Rez=1 =11.06x109. NuHeat transfer for plate length L=0.25 m. Q2-0.25 = HZ-0.25A, (T) –T.) Q2-0.25 = 133.404x(0.25x1)(350 – 300) QL=0.25 = 1667.55